1.What pressure, in mmHg, is required to contain 0.023 moles of nitrogen gas in a 4.2 L container at a temperature of 20.°C?
1
Expert's answer
2019-05-10T08:04:49-0400

To count out the required pressure, we will use this formula:

PV = nRT

As R=82.057 cm3× atm / (mol × K), then we must convert l into cm3 and temperature into K.

V = 4.2 l = 4200 cm3

T = t + 273.15 = 20 + 273.15 = 293.15 K

P = nRT/V = 0.023 × 82.057 × 293.15 / 4200 = 0.1317 (atm) = 100.092 = mmHg.